# RadarLLM: Adapting Pretrained Large Language Models for Marine Radar Target Detection with Preference-aware Loss

### v1

@@ Abstract
Recent advances in pre-trained large language models (LLMs) have demonstrated their capacities to capture universal knowledge, making them promising general-purpose optimization solvers for wireless signal processing. Motivated by these findings, we take the first step towards fine-tuning pre-trained LLMs for the effective analysis of radar signal features in marine target detection tasks. Nevertheless, directly fine-tuning pre-trained LLMs on marine target detection tasks tends to suffer from pronounced overfitting, particularly in challenging low signal-to-clutter ratio (SCR) scenarios. This overfitting primarily stems from the model’s tendency to memorize spurious or noisy feature patterns rather than learning discriminative structures that generalize well to unseen data. To address this challenge, we introduce RadarLLM, a novel fine-tuning framework that utilizes an effective preference-aware loss. Unlike conventional training strategies that uniformly optimize all feature tokens, this loss function selectively optimizes different feature patches based on their online evaluated learning values, thus guiding the model to focus on the most generalizable patterns during optimization. We theoretically demonstrate the effectiveness of the evaluated learning values by transforming the problem as selecting useful feature tokens. Extensive experiments on real-world marine radar datasets show that 1) the proposed loss function is much better than the original one, with particularly significant gains in challenging low SCR scenarios and 2) RadarLLM consistently outperforms state-of-the-art baselines across diverse detection scenarios, with particularly notable gains under limited training data conditions.
@@ I Introduction
To address the aforementioned challenges, this paper introduces RadarLLM, an innovative framework for learning from radar echo signals. First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens. Second, we train a small reference model based on the validation samples and calculate the loss for each feature token of the training samples to estimate its generalization potential. Third, we construct a target detector based on a pre-trained LLM, then fine-tune the model based on low-rank adaptation (LoRA) [26] and a novel preference-aware loss function, which is inspired by the work in NLP [27] and CV [28]. This loss introduces a selective training strategy by assigning training weights to different feature patches according to their learning values, which greatly improving model performance in low SCR environments. Finally, we retrain a classification head to enhance detection performance and control the false alarm rate.
@@
    We propose RadarLLM, a novel framework combining a lightweight reference model and a pre-trained LLM for marine radar target detection. To the best of our knowledge, we are the first to show that well-known LLM can be a strong starting point for intelligent radar signal processing.
@@
    For fine-tuning of the framework, we propose a novel preference-aware loss function which greatly improves RadarLLM’s performance in a low SCR environment. RadarLLM can selectively optimize on useful feature tokens with high learning values while reducing the risk of overfitting.
@@
    Our extensive experiments confirm the effectiveness of RadarLLM. It attains new state-of-the-art detection performance under sufficient training data, and constrained training data settings on popular marine radar datasets.
@@ II-A Deep Learning Powered Marine Radar Target Detection
Deep learning models with elaborately crafted architectures have demonstrated great promise in marine target detection. Among them, Chen et al. [7] design a dual-channel CNN (DCCNN)-based structure detector that extracts both amplitude and time-frequency information from signals to achieve target detection. Qu et al. [8] introduce a CNN architecture augmented with asymmetric convolution layers to capture deep features of the time-frequency distribution. Xu et al. [9] propose a target detector in the case of limited training samples, utilizing pre-trained CNN to extract features from time-frequency spectrogram. Wan et al. [14] propose a sequence-based target detection framework that leverages instantaneous phase, Doppler spectrum, and short-time Fourier transform features, in conjunction with a bidirectional long short-term memory (Bi-LSTM) network. Wang et al. [11] propose a target detector based on complex value U-Net (CV-UNet), which realizes clutter suppression based on the amplitude-phase characteristics of radar echoes, and then achieves target detection. Su et al. [17] introduce a graph neural network (GNN) approach for radar target detection, which constructs spatio-temporal adjacency matrices, extracts hierarchical features through convolutional graph operations, and outputs detection results via nonlinear vertex embeddings.
In addition to traditional supervised learning, semi-supervised, unsupervised, and incremental learning approaches are proposed to further advance radar target detection. Wang et al. [29] propose a self-evolving framework for maritime radar target detection using semi-supervised learning, which enhances detection performance through unsupervised sample selection, data augmentation, and model optimization. Xia et al. [10] introduce an unsupervised contrastive learning framework that learns discriminative representations between targets and clutter from unlabeled data. Wang et al. [30] propose an incremental learning-based target detection method, which can continuously adapt the NN model in real time according to environmental changes. These methods provide compelling evidence that deep learning can significantly enhance detection performance without the need for explicit mathematical modeling. However, it is important to recognize certain limitations inherent in these methods. Current methods rely on relatively small NN models, which possess limited feature recognition capabilities and are highly sensitive to environmental variations.
@@ II-B Language Model Powered Signal Processing Task
Recently, LLMs such as ChatGPT and DeepSeek have demonstrated exceptional capabilities in natural language understanding, code generation, math problem solving, and human-like text generation [31]. Some studies also demonstrate their potential in the signal processing (SP) task. For example, Zhou et al. [18] propose a unified time series analysis framework by fine-tuning frozen GPT2 [32] and achieving SOTA performance on various datasets. Liu et al. [24] propose a GPT2-empowered channel prediction framework to improve prediction accuracy. Sheng et al. [33] utilize the GPT2 model to develop an effective and robust beam prediction method. Zheng and Dai [21] propose a multi-task LLM framework to satisfy the requirements of different wireless communication tasks. IOT-LLM and Penetrative AI [34, 35] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks. However, recent studies [36] have raised concerns about the universal applicability of LLM in signal processing tasks. LLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks. In marine radar target detection under low SCR conditions, directly fine-tuning a pre-trained LLM often leads to severe overfitting issues, limiting its parameter transfer capability. We counteract this shortfall with a preference-aware loss through selectively training on useful feature patches.
@@ III-A Overview of the RadarLLM Framework
In this study, we present RadarLLM, a LLM–empowered framework for marine radar target detection. The framework introduces a selective patch-level fine-tuning strategy that counteracts the severe overfitting problem caused by the low SCR environment. The overall pipeline of RadarLLM is illustrated in Fig. 1 and shown in Algorithm 1.
First, five complementary sequence features are extracted from radar echoes and partitioned into uniform patches to form a multi-patch representation. Next, a lightweight reference model rapidly scores each patch to quantify its generalization potential. Then, we fine-tune a pre-trained LLM backbone via a preference-aware loss and Low-Rank Adaptation (LoRA). Finally, an autoencoder-based classification head is re-trained to enhance detection performance and control the false alarm rate. The detailed design and rationale of each module are elaborated in the following subsections.
Algorithm 1 RadarLLM
@@
III-B Stage 1: Sequence Feature Extraction
@@
As shown in Fig. 2, the five sequence features reveal clear distinctions between target and clutter samples.
@@ III-C Structure of the RadarLLM Network
Initially, for a batch of $B$ samples, the input feature matrix is constructed by concatenating the IP feature $F_{\mathrm{IP}}$, DSE feature $F_{\mathrm{DSE}}$, SMS feature $F_{\mathrm{SMS}}$, amplitude feature $F_{\mathrm{Amp}}$, and DP features $F_{\mathrm{DP}}$.
To adapt the feature patches to the input format required by the pre-trained language model, we employ a fully connected layer to project ${F}^{\mathrm{P}}$ into an embedded representation ${F}^{\mathrm{EB}}\in\mathbb{R}^{B\times K\times L^{\prime}}$. This is followed by the addition of a positional encoding ${E}^{\mathrm{pos}}$, which incorporates either relative or absolute positional information into the embedded features:
${E}^{\mathrm{pos}}_{b,k,2l}=\sin(\frac{k}{10000^{2l/L^{\prime}}}),\ {E}^{\mathrm{pos}}_{b,k,2l+1}=\cos(\frac{k}{10000^{2l/L^{\prime}}}),$ (11)
where $b$ represents the $b$-th sample, $k$ represents the position index of the token, and $l$ denotes the feature dimension index. The positional encoding ${E}^{\mathrm{pos}}$ is added to the embedding ${F}^{\mathrm{EB}}$ to produce a enriched embedding:
${F}^{\mathrm{PE}}={F}^{\mathrm{EB}}+{E}^{\mathrm{pos}}.$ (12)
The enriched embedding ${F}^{\mathrm{PE}}$ is passed through initial several layers of the GPT2 encoder for further feature extraction. To efficiently fine-tune the backbone model, we apply Low-Rank Adaptation (LoRA) [26] to the query ($\mathbf{W}_{Q}$) and value matrices ($\mathbf{W}_{V}$) in the multi-head attention layers and directly fine-tune the Layer Normalization (LayerNorm) parameters. The LoRA fine-tuning process is illustrated in Fig. 4. This strategy introduces lightweight, trainable low-rank matrices to the fixed pre-trained weights, substantially reducing the computational and memory requirements compared to full parameter fine-tuning. Formally, the adapted model generates rich contextual feature representations as follows:
${F}^{\mathrm{LLM}}=\operatorname{LLM}_{\mathrm{LoRA}}\left({F}^{\mathrm{PE}}\right)\in\mathbb{R}^{B\times K\times L^{\prime}},$ (13)
where $\operatorname{LLM}_{\mathrm{LoRA}}(\cdot)$ represents the backbone network of the language model encoder enhanced with LoRA modules. To stabilize the training process, a Layer Normalization (LayerNorm) operation is applied to the output representation ${F}^{\mathrm{LLM}}$.
$F^{\mathrm{LN}}=\mathrm{LayerNorm}({F}^{\mathrm{LLM}}),$ (14)
where $\mathrm{LayerNorm}(\cdot)$ is defined element-wise for each position $k$ and feature dimension $l$ as:
$F_{b,k,l}^{\mathrm{LN}}=\gamma_{b,l}\cdot\frac{F_{b,k,l}^{\mathrm{LLM}}-\mu_{b,l}}{\sqrt{\sigma_{b,l}^{2}+\epsilon}}+\beta_{b,l},$ (15)
where $\mu_{b,l}$ and $\sigma_{b,l}^{2}$ is computed as:
$\mu_{b,l}=\frac{1}{K}\sum_{k=1}^{K}F_{b,k,l}^{\mathrm{LLM}},\quad\sigma_{b,l}^{2}=\frac{1}{K}\sum_{k=1}^{K}\left(F_{b,k,l}^{\mathrm{LLM}}-\mu_{b,l}\right)^{2}.$ (16)
Here, $\gamma_{b,l}$ and $\beta_{b,l}$ are learnable parameters, and $\epsilon$ is a small constant for numerical stability. $F^{\mathrm{LN}}$are then passed through two FC layers to obtain the final output:
$F^{\mathrm{fi}}=FC(FC(F^{\mathrm{LLM}}))\in\mathbb{R}^{B\times K\times 2}.$ (17)
@@
III-D Stage 2: Compute Loss for Each Token Based on the Reference Model
The architecture of the reference model is illustrated in Fig. 5. As described in Section III-C, the input patches are first projected into a latent Transformer space of dimension $D$ and augmented with learnable additive positional embeddings. This process yields the enriched representation $F^{\mathrm{PE}^{\prime}}\in\mathbb{R}^{B\times K\times D}$, which is subsequently processed by a Transformer encoder. Within the encoder, the multi-head attention (MHA) mechanism functions as follows: the input $F^{\mathrm{PE}^{\prime}}$ is linearly projected into $H$ independent attention heads. For each head $h$, query, key, and value matrices $\mathbf{W}_{Q_{h}}\in\mathbb{R}^{B\times D\times d_{k}}$ $\mathbf{W}_{K}\in\mathbb{R}^{B\times D\times d_{k}}$ $\mathbf{W}_{V}\in\mathbb{R}^{B\times D\times d_{v}}$ are used to compute:
$\mathbf{Q}_{h}=F^{\mathrm{PE}^{\prime}}\mathbf{W}_{Q_{h}},\quad\mathbf{K}_{h}=F^{\mathrm{PE}^{\prime}}\mathbf{W}_{K_{h}},\quad\mathbf{V}_{h}=F^{\mathrm{PE}^{\prime}}\mathbf{W}_{V_{h}},$ (18)
where $d_{k}=d_{v}=D/H$. The attention output of each head is then computed as:
$\mathbf{O}_{h}=\operatorname{Softmax}\left(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}{}^{\bm{\top}}}{\sqrt{d_{k}}}\right)\mathbf{V}_{h}.$ (19)
Outputs from all heads are concatenated along the feature dimension and projected back to dimension $D$. The MHA block further incorporates Batch Normalization layers and a feed-forward network (FFN) equipped with residual connections, resulting in the final encoded representation $F^{\mathrm{TF}}\in\mathbb{R}^{B\times K\times D}$. Subsequently, LayerNorm is applied, followed by two fully connected layers to yield the final output $F^{\mathrm{fi}^{\prime}}\in\mathbb{R}^{B\times K\times 2}$.
Figure 5: Overview of the reference model.
III-E Stage 3: Fine-tune RadarLLM Network Based on the Preference-aware Loss
III-E1 Methodology
To improve fine-tuning of LLMs for radar target detection in low SCR scenarios, we propose a token-level preference-aware loss function. In such conditions, many feature patches are either severely corrupted by noise or trivially easy, offering limited value for effective learning. Uniformly optimizing all patches risks overfitting by forcing the model to learn from uninformative or misleading inputs. To address this, we introduce a dynamic weighting mechanism that prioritizes tokens with higher learning value, those expected to contribute more to generalization. Specifically, our approach utilizes a dual-model framework consisting of a lightweight reference model $\theta_{r}$ and a high-capacity target model $\theta_{t}$, which leverages LLMs.
@@ III-E1 Methodology
  • Target Model: A significantly larger model built upon a pre-trained LLM, parameterized by $\theta_{t}$, as described in Section III-C.
@@
III-E2 Theoretical Insight of the Effectiveness of Feature Token Importance

### v2

@@ Abstract
Recent advances in pre-trained large language models (LLMs) have demonstrated their capacities to capture universal knowledge, making them promising general-purpose optimization solvers for wireless signal processing. Motivated by these findings, we take the first step towards fine-tuning pre-trained LLMs for the effective analysis of radar signal features in marine target detection tasks. Nevertheless, directly fine-tuning pre-trained LLMs on marine target detection tasks tends to suffer from pronounced overfitting, particularly in challenging low signal-to-clutter ratio (SCR) scenarios. This overfitting is mainly due to the model’s tendency to memorize noisy feature patterns rather than learning discriminative structures that generalize well to unseen data. To address this challenge, we introduce RadarLLM, a novel fine-tuning framework that utilizes an effective preference-aware loss. Unlike conventional training strategies that uniformly optimize all feature tokens, this loss function selectively optimizes different feature patches based on their online evaluated learning values, thus guiding the model to focus on the most generalizable patterns during optimization. We theoretically demonstrate the effectiveness of the evaluated learning values by transforming the problem as selecting useful feature tokens. Extensive experiments on real-world marine radar datasets show that 1) the proposed loss function outperforms the original one, showing particularly significant improvements under challenging low SCR conditions, with an average performance gain of 9.9% and 2) RadarLLM consistently outperforms state-of-the-art baselines in diverse detection scenarios, with particularly notable gains under limited training data conditions.
@@ I Introduction
To address the aforementioned challenges, this paper introduces RadarLLM, an innovative framework designed for learning from radar echo signals. First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens. Second, we train a small reference model based on the validation samples and calculate the loss for each feature token of the training samples to estimate its generalization potential. Then, we construct a target detector based on a pre-trained LLM and fine-tune the model using a lightweight finetuning module, along with a novel preference-aware loss function inspired by the work in NLP [26] and CV [27]. This loss function introduces a selective training strategy by assigning different training weights to feature patches according to their online evaluated learning values, significantly improving model performance in low SCR environments. Finally, we retrain the classification head to enhance detection performance and control the false alarm rate.
@@
    We propose RadarLLM, a novel framework combining a lightweight reference model and a pre-trained LLM for marine radar target detection. To the best of our knowledge, we are the first to show that the well-known LLMs can be a strong starting point for intelligent radar signal processing.
@@
    For the fine-tuning of the LLMs, we propose a novel preference-aware loss function. This loss function can selectively optimize useful feature tokens with high learning values, significantly reducing the risk of model overfitting.
@@
    Our extensive experiments confirm the effectiveness of RadarLLM. It attains new state-of-the-art detection performance under sufficient training data and constrained training data settings on popular marine radar datasets.
@@ II-A Deep Learning Powered Marine Radar Target Detection
Deep learning models with elaborately crafted architectures have demonstrated great promise in marine target detection. Among them, Chen et al. [7] design a dual-channel CNN (DCCNN)-based structure detector that extracts both amplitude and time-frequency information from signals to achieve target detection. Qu et al. [8] introduce a CNN architecture augmented with asymmetric convolution layers to capture deep features of the time-frequency distribution. Xu et al. [9] propose a target detector in the case of limited training samples, utilizing pre-trained CNN to extract features from time-frequency spectrogram. Wan et al. [14] propose a sequence-based target detection framework that leverages instantaneous phase, Doppler spectrum, and short-time Fourier transform features, in conjunction with a bidirectional long-short-term memory (Bi-LSTM) network. Wang et al. [11] propose a target detector based on complex value U-Net (CV-UNet), which performs clutter suppression based on the amplitude-phase characteristics of radar echoes, and then achieves target detection. Su et al. [17] introduce a graph neural network (GNN) approach for radar target detection, which constructs spatio-temporal adjacency matrices, extracts hierarchical features through convolutional graph operations, and outputs detection results via nonlinear vertex embeddings.
In addition to traditional supervised learning, semi-supervised, unsupervised, and incremental learning approaches are proposed to further advance radar target detection. Wang et al. [28] propose a self-evolving framework for maritime radar target detection using semi-supervised learning, which enhances detection performance through unsupervised sample selection, data augmentation, and model optimization. Xia et al. [10] introduce an unsupervised contrastive learning framework that learns discriminative representations between targets and clutter from unlabeled data. Wang et al. [29] propose an incremental learning-based target detection method, which can continuously adapt the NN model in real time according to environmental changes. These methods provide compelling evidence that deep learning can significantly enhance detection performance without the need for explicit mathematical modeling. Despite their initial promise, these methods are inherently constrained by their dependence on relatively compact neural networks. Such models exhibit limited representational capacity, which hinders their ability to learn complex, high-dimensional feature representations. Consequently, they are particularly susceptible to performance degradation in the face of diverse and nonstationary environmental conditions. These challenges motivate us to first explore the use of LLMs for radar target detection.
@@ II-B Language Model Powered Signal Processing Task
Recently, LLMs such as ChatGPT and DeepSeek have demonstrated exceptional capabilities in natural language understanding, code generation, math problem solving, and human-like text generation [30]. Some studies also demonstrate their potential in the signal processing (SP) task. For example, Zhou et al. [18] propose a unified time series analysis framework by fine-tuning frozen GPT2 [31] and achieving SOTA performance on various datasets. Liu et al. [24] propose a GPT2-empowered channel prediction framework to improve prediction accuracy. Sheng et al. [32] utilize the GPT2 model to develop an effective and robust beam prediction method. Zheng and Dai [21] propose a multi-task LLM framework to satisfy the requirements of different wireless communication tasks. IOT-LLM and Penetrative AI [33, 34] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks. However, recent studies [35] have raised concerns about the universal applicability of LLM in signal processing tasks. LLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks. In marine radar target detection under low SCR conditions, directly fine-tuning a pre-trained LLM often leads to severe overfitting issues, limiting its parameter transfer capability. To mitigate this limitation, we introduce a preference-aware loss that selectively trains on informative feature patches.
@@ III-A Overview of the RadarLLM Framework
In this study, we present RadarLLM, an LLM–empowered framework tailored for the marine radar target detection task. RadarLLM training protocol is illustrated in Fig. 1 and shown in Algorithm 1.
To enable lightweight yet effective fine-tuning of LLMs for radar target detection, the proposed RadarLLM framework integrates four tightly coupled modules into a unified optimization pipeline. First, the sequence feature extraction and patching module (Section III-B) derives five discriminative sequence features from the multi-domain transformation of radar echo signals. Then a patching operation is applied to capture local semantic patterns while reducing computational complexity. Second, a lightweight finetuning module (Section III-C) is introduced to efficiently adapt the LLM backbone to the radar task, preserving its universal knowledge contained while minimizing parameter updates. Third, a preference-aware loss function (Section III-D), which utilizes a lightweight reference model, enables selective training on informative feature patches and effectively alleviates overfitting problem. Finally, an autoencoder-based classification head (Section III-E) is retrained to enhance the detection rate and maintain a controllable false alarm rate.
As illustrated in Fig. 1, the LLM-based model first employs an input embedding layer to project the feature patches into the same latent space as the pre-trained LLM. The embedded representations are then processed by the initial encoder layers of the pre-trained LLM to extract more discriminative features. Subsequently, a layer normalization operation refines these features, yielding the normalized representation $F^{\mathrm{LN}}$, which enhances consistency and stability. Finally, an output linear layer generates token-level detection results. For the structure of the reference model, a standard Transformer encoder [36] is utilized to process the input feature patches and produce token-level detection outputs, as illustrated in Fig. 2.
Algorithm 1 RadarLLM Training Protocol
@@
Figure 2: Overview of the reference model.
III-B Sequence Feature Extraction and Patching
@@
As illustrated in Fig. 3, the five extracted sequence features exhibit pronounced discriminability between target and clutter samples. To better structure the input for downstream processing, the extracted features are partitioned into non-overlapping patches. This patching approach offers several key advantages: it helps retain local semantic information, reduces the computational and memory burden of attention mechanisms. For a mini-batch of $B$ samples, the input feature matrix is constructed by concatenating the IP feature $F_{\mathrm{IP}}$, DSE feature $F_{\mathrm{DSE}}$, SMS feature $F_{\mathrm{SMS}}$, amplitude feature $F_{\mathrm{Amp}}$, and DP features $F_{\mathrm{DP}}$.
@@
III-C Lightweight Fine-tuning Module for the LLM Backbone
Inspired by the strategy proposed in [18], we adopt a lightweight fine-tuning module that jointly updates the layer normalization parameters and integrates a Low-Rank Adaptation (LoRA) module [37]. LoRA is a parameter-efficient adaptation technique that freezes the pre-trained model weights while injecting trainable low-rank decomposition matrices into selected layers.
This design offers two key advantages. First, given the immense parameter scale of LLMs, full fine-tuning is often computationally prohibitive. This module mitigates this limitation by updating only compact low-rank matrices, thereby reducing trainable parameters by several orders of magnitude. Second, this module alleviates catastrophic forgetting of pre-trained knowledge during adaptation, as the newly acquired information is encoded within a subspace of lower rank than the original representation. Consequently, this module enables the finetuned model to retain its pre-trained general-purpose knowledge while efficiently adapting to downstream radar tasks.
Concretely, we fine-tune the query and value matrices, $\mathbf{W}_{Q}$ and $\mathbf{W}_{V}$, within each attention block of the LLM backbone, as illustrated in Fig. 5. LoRA introduces two trainable rank decomposition matrices, $A$ and $B$, attached to the frozen pre-trained weight matrix $\mathbf{W}$. The modified transformation is expressed as:
$\mathbf{W}\leftarrow\mathbf{W}+AB^{T},$ (11)
where $\mathbf{W}\in\mathbb{R}^{d{1}\times d{2}}$, $A\in\mathbb{R}^{d_{1}\times r}$, $B\in\mathbb{R}^{d_{2}\times r}$, and $r\ll\min(d_{1},d_{2})$. The matrices are initialized as $A\sim\mathcal{N}(0,\sigma^{2})$ and $B=0$. During fine-tuning, $\mathbf{W}$ remains frozen while $A$ and $B$ are optimized. Owing to the low-rank constraint ($r\ll\min(d_{1},d_{2})$), this approach dramatically reduces the number of trainable parameters compared to full-parameter fine-tuning.
@@ III-C Lightweight Fine-tuning Module for the LLM Backbone
III-D Preference-aware Loss Function
III-D1 Methodology
To enhance the fine-tuning of LLMs for radar target detection under low SCR conditions, we introduce a token-level preference-aware loss function. In such scenarios, a substantial proportion of feature patches contribute little information for effective learning. Uniform optimization over all feature patches risks overfitting by coercing the model to memorize uninformative or even misleading patterns. To mitigate this, we propose a dynamic re-weighting scheme that adaptively assigns higher learning weights to feature patches exhibiting superior generalization potential. Specifically, our approach uses a dual-model framework consisting of a lightweight reference model $\theta_{r}$ and a high-capacity target model $\theta_{t}$, which leverages LLMs.
@@ III-D1 Methodology
  • Target Model: A significantly larger model built upon a pre-trained LLM, parameterized by $\theta_{t}$.
@@
III-D2 Theoretical Insight of the Effectiveness of Feature Token Importance